### v2

# Distributed Rhombus Formation of Sliding Squares

Abstract
The sliding square model is a widely used abstraction for studying self-reconfigurable robotic systems, where modules are square-shaped robots that move by sliding or rotating over one another. In this paper, we propose a novel distributed algorithm that allows a group of modules to reconfigure into a rhombus shape, starting from an arbitrary side-connected configuration. It is connectivity-preserving and operates under minimal assumptions: one leader module, common chirality, constant memory per module, and visibility and communication restricted to immediate neighbors. Unlike prior work, which relaxes the original sliding square move-set, our approach uses the unmodified move-set, addressing the additional challenge of handling locked configurations. Our algorithm is sequential in nature and operates with a worst-case time complexity of $\mathcal{O}(n^{2})$ rounds, which is optimal for sequential algorithms. To improve runtime, we introduce two parallel variants of the algorithm. Both rely on a spanning tree data structure, allowing modules to make decisions based on local connectivity. Our experimental results show a significant speedup for the first variant, and linear average runtime for the second variant, which is worst-case optimal for parallel algorithms.
keywords:
modular robots, distributed systems, sliding squares
category:
[…]
1 Introduction
Edge-connected configurations of square modules, which can reconfigure by sliding over or rotating around one another (see LABEL:subfig:def-moves), are a well-established theoretical model for self-reconfigurable robotic systems. There are numerous applications such as search-and-rescue missions or autonomous construction, where modules collectively navigate complex terrains or form shapes while preserving connectivity. Reconfiguration planning in the sliding square model has primarily focused on centralized approaches. In contrast, we propose a fully distributed algorithm that enables modules to autonomously self-organize into a rhombus shape under minimal assumptions about their initial positions, computational capabilities, and sensory range. The rhombus serves as an efficient intermediate shape due to its compactness, minimal bounding box, and lack of holes.
1.1 Our Contribution
We introduce a distributed reconfiguration algorithm for the sliding square model that transforms any initially side-connected configuration with a designated leader and common chirality into a rhombus shape. These two assumptions are minimal but necessary to break initial symmetries. The algorithm has a worst-case runtime of $\mathcal{O}(n^{2})$ rounds, which is optimal among solutions that build up any shape sequentially. It operates under strict local communication and visibility constraints, requires only constant memory per module, and maintains connectivity throughout execution. In contrast to previous distributed solutions, which relax the original sliding square move-set, our algorithm utilizes the unmodified move-set, introducing the additional challenge of handling locked configurations. While our main focus is the rhombus, the algorithm generalizes to any shape that can be sequentially constructed, provided that at every step, any two cells in the shape remain connected by a shortest path within the shape.
To improve efficiency, we introduce two parallel variants of the algorithm. Both variants rely on an auxiliary spanning tree structure maintained throughout execution, allowing modules to detect certain moves that preserve connectivity. The first variant permits modules to move only if such moves improve their depth relative to the spanning tree (without explicitly computing or storing depth), resulting in a significant speedup over the sequential version. However, some configurations remain locked, limiting parallelization and requiring sequential progression. The second variant allows modules to move whenever possible, showing an average-case linear runtime, as demonstrated by our experiments.
1.2 Related Work
Reconfiguration in the sliding square model, first introduced by Fitch et al. [15], typically involves transforming configurations into a canonical intermediate shape before reaching a desired target shape. Dumitrescu and Pach [12] were the first to provide a universal 2D reconfiguration algorithm in this model, achieving a worst-case optimal runtime of $\mathcal{O}(n^{2})$ moves using a straight line as the canonical shape. Moreno and Sacristán [27] adapted this into an in-place version, restricting all intermediate states to lie within the 1-offset bounding box of the initial and target configurations, while still requiring $\Theta(n^{2})$ moves. Akitaya et al. [2] presented the first input sensitive algorithm, achieving $\mathcal{O}(nP)$ complexity, where $P$ is the maximum perimeter of the bounding boxes, using $xy$-monotone shapes and proving that minimizing the number of moves is NP-hard. Abel et al. [3] give the first universal reconfiguration algorithm for sliding cubes in three and higher dimensions, achieving an optimal $\mathcal{O}(n^{2})$ bound and an in-place variant where all but one module stay within the bounding boxes of the start and target configurations. Kostitsyna et al. [26] achieve asymptotically optimal in-place reconfiguration for sliding squares, with the number of moves proportional to the total coordinate sum over all cubes.
All aforementioned algorithms are centralized and sequential. The first distributed and synchronous approach was by Dumitrescu et al. [11], who reconfigure $y$-monotone shapes into a line, assuming global communication. Hurtado et al. [21] later gave the first distributed universal algorithm with optimal $\mathcal{O}(n)$ rounds, but with a relaxed move set, a leader requiring $\mathcal{O}(n)$ memory, and $\mathcal{O}(\log n)$ memory in other modules (despite their claim of constant memory, as storing integers up to $n$ requires $\log n$ bits). They further assume a vision and communication range of $2$. Wolters [31] presented parallel versions for reconfiguring $xy$-monotone shapes, though their work lacks correctness proofs. Recently, Akitaya et al. [5] gave the first centralized theoretical results for parallel reconfiguration in the sliding square model, achieving worst-case optimal makespan $\mathcal{O}(P)$ and showing NP-completeness for constant-makespan decisions.
In the limited sliding cube model, modules are restricted to sliding moves only, and reconfiguration is achieved through meta-modules, small groups of modules that act as single units. Kawano proposed a series of increasingly efficient algorithms in this model, culminating in a linear-time solution [23, 24, 25].
Pivoting models, which allow only corner rotations, are significantly harder. Akitaya et al. [1, 6] demonstrated PSPACE-hardness for several pivoting reconfiguration problems in both 2D and 3D. A notable workaround uses five auxiliary modules (called musketeers) to achieve universal reconfiguration in $\mathcal{O}(n^{2})$ pivot moves [4]. Non-universal strategies focus on configurations excluding certain forbidden substructures. Sung et al. [29] identified three such subconfigurations, allowing reconfiguration in $\mathcal{O}(n^{2})$ moves, which was extended by Feshbach and Sung [14] to include non-convex holes.
Other relevant problems include coordinated motion planning, where multiple agents navigate a shared space while maintaining connectivity [13, 30], nest-building inspired by biological systems to form compact structures [7], and bounding-box construction [28].
2 A Distributed Rhombus Formation Algorithm
In this section, we describe our distributed rhombus formation algorithm.
3 Analysis
In the following, we prove that our distributed rhombus formation algorithm reconfigures any initially side-connected configuration $(M_{0},s_{0},l)$ with a leader in cell $l\in M_{0}$ into a rhombus centered at $l$, with its outermost layer possibly partially filled.
4 Utilizing Parallelization
In the previous sections, we presented and analyzed a distributed rhombus formation algorithm in which modules fill positions within the rhombus sequentially in $\mathcal{O}(n^{2})$ rounds.
5 Parallel Variants and Experimental Evaluation
To further improve performance, we developed two variants, V1 and V2, that increase the degree of parallelism. V1 allows leaf modules to perform state swaps similar to those of the Active module whenever their movement trajectories intersect. This enables progress in configurations like the one in LABEL:subfig:locked-simple without waiting for requests. Combined with the parallelization strategy from Section˜4, V1 already yields a significant speedup over the purely sequential algorithm in most cases. Importantly, the correctness of the algorithm does not depend on the specific movement choices of the leaves, as long as leaves are not allowed to attach to the request path. However, since leaf modules are restricted to moves that improve their depth in $\mathcal{T}{}$, certain configurations still become locked, in which case the runtime is asymptotically not better than the sequential algorithm.
V2 addresses this by relaxing the depth-improvement requirement. Leaves move as in V1, with one addition: a leaf proceeds with an available move even if it cannot locally verify that the move improves its depth in $\mathcal{T}{}$. Such moves may temporarily increase a leaf’s depth and delay its progress, but they can also turn its parent into a leaf, unlocking further parallel movement. To prevent unproductive cycles in which leaves repeatedly step over each other, we do not allow leaves to attach themselves to another Passive leaf. In our experiments, V2 consistently outperforms V1 and appears to achieve linear runtime on average.
Simulation Setup and Evaluation
We evaluated the performance of the sequential algorithm and both parallel variants (V1 and V2) on structured and random configurations using a custom-built simulator. For each configuration type and size, we ran multiple independent simulations.
The structured input configurations considered are lines, where modules are arranged in a straight horizontal segment; spirals, generated by placing modules in a square spiral pattern; and chains, constructed as cascaded sequences of critical pairs. In all structured configurations, the initial leader is explicitly placed at a degree-one module. The top row in Figure˜1 shows the number of rounds required to reconfigure a line and a chain into a rhombus, for configurations of size 10 to 120 in increments of 10. Spiral configurations are excluded from the plots since they follow the same trend as line configurations. We also include the Earth Mover’s Distance (EMD) as a lower bound for any sequential solution, and the number of modules $n$ as a lower bound for V1 and V2.
Figure 1: Comparison of the runtime of the sequential rhombus formation algorithm (Seq) with its two parallel variants (V1 and V2). The top row shows the runtime on line and chain configurations; the second row shows the runtime on random percentage configurations. In all but the top row, the right plot displays the same data but excludes the sequential runtime for better resolution. The third and fourth rows compare runtime against the Earth Mover’s Distance (EMD) and the leader’s eccentricity for random configurations (both percentage-based and Perlin noise-based) of fixed size $n=200$. Scatter points indicate individual runs, while lines are smoothed using LOWESS (locally weighted regression over nearby points).
As expected, the runtime of the sequential algorithm grows quadratically on both line and chain configurations, which are worst-case inputs in the sequential setting. In contrast, the runtimes of V1 and V2 coincide on line and spiral configurations, where leaves are not blocked during reconfiguration. However, for chain configurations, V1 yields only a constant-factor improvement over the sequential algorithm, while V2 appears to achieve linear runtime.
We used two different types of random input configurations:
  1. 1.
    Percentage-based Configurations: Starting from a filled rectangle, modules were removed uniformly at random while preserving connectivity, until a specified percentage remained.
  2. 2.
    Growth-based Configurations (Perlin Noise): These configurations were grown from a single seed module. In each step, all empty cells side-adjacent to existing modules were assigned a weight based on a fixed 2D Perlin noise map (a type of gradient noise). Among the 10 cells with the highest weights, one was selected uniformly at random and added to the configuration. This method produces irregular, organic shapes with many holes.
For each rectangle of dimension $(x,y)$ with $x$ ranging from 6 to 20 in steps of 2, and $y$ from $x$ to 20 in steps of 2, we generated 10 configurations with 10%, 30%, 50%, 70%, and 90% of the original modules retained, resulting in a total of 1800 simulations. The second row in Figure˜1 shows the results of all those runs.
Both variants significantly improve runtime over the purely sequential algorithm and even outperform the theoretical lower bound given by the Earth Mover’s Distance (EMD). This is expected, as the parallel algorithm allows multiple modules to move simultaneously, while the EMD assumes only one module moves per round. Additionally, V2 consistently outperforms V1 across all simulations. In conclusion, the speedup from increased parallelism in V2 outweighs the benefit of shorter paths in V1.
To analyze which configuration parameters influence runtime, we generated over 1000 random configurations, both percentage-based and growth-based, each with 200 modules. For each configuration, we computed several metrics: its diameter, the number of holes and pseudo-holes, the leader module’s eccentricity (i.e., the largest distance to any non-leader module), its closeness (i.e., the mean eccentricity of all modules), the EMD towards the rhombus, and the perimeter (i.e., the number of edges between empty and occupied cells).
Analyzing the plots, we find that only the leader’s eccentricity, closeness, and EMD correlate with runtime. As eccentricity and closeness show nearly identical trends and other parameters appear uncorrelated, we include only the plots for eccentricity and EMD, shown in the third and fourth rows of Figure˜1, respectively.
The runtime increases roughly linearly with EMD, as expected: larger differences between the initial and target configurations require more total module movement. In contrast, the runtime grows sublinearly, approximately following a square-root curve, with the leader’s eccentricity. This is likely because distant modules rarely move directly to the target cell. Instead, they often fill critical pairs along the way, enabling closer modules to move first. As a result, the rhombus grows, reducing the distance that distant modules must traverse in later rounds. In the EMD plot, V2 shows a smooth, steady linear runtime with minimal outliers, while V1 shows higher variance. This suggests that locked leaves, as constructed manually in earlier examples, frequently occur even in random configurations, causing delays under V1 but not under V2.
6 Conclusion
We presented a distributed reconfiguration algorithm for the sliding square model that transforms side-connected initial configurations into a rhombus. The algorithm operates under strict local communication and constant memory constraints, while maintaining connectivity throughout execution. It runs sequentially with a worst-case runtime of $\mathcal{O}(n^{2})$, which is optimal for sequential solutions under these constraints. While our focus was on the rhombus, the algorithm generalizes to any target shape that is sequentially constructible and weakly convex at each construction step, including, for example, lines and filled rectangles.
To improve runtime, we introduced two parallel variants V1 and V2. Simulations show that both variants outperform the sequential algorithm and even surpass the theoretical lower bound for sequential runtimes given by the Earth Mover’s Distance, with V2 achieving the greater speedup. We identified key factors influencing runtime: it scales linearly with EMD and follows a square-root curve with respect to the leader’s eccentricity and closeness. V2, in particular, showed a smooth runtime curve with low variance, suggesting it can handle configurations that tend to become locked under V1.
Future work includes extending the algorithm to non-weakly convex shapes, exploring transformations from the rhombus to arbitrary configurations, and explore lower bounds to clarify the complexity of distributed reconfiguration in the sliding square model.